Litego school of biology student used a microscope with ×40 objective lens and ×15 eye piece lens which had a 2mm radius. Calculate the area of the field of view in micrometers
1
Expert's answer
2020-04-30T15:45:57-0400

Field view = Field Number (FN) / Objective magnification

x40 * x15 / 2=2/800 =0.0025

1 mm=1000 micrometres

0.0025*100=2.5 micrometres.